COMPOSITE APPROACHES FOR AB INITIO SPECTROSCOPY: THE CCN, CCSb, AND HNNO RADI-CALS

\underline{KIRK A. PETERSON}, J. GRANT HILL, JAMES SHEAROUSE, Department of Chemistry, Washington State University, Pullman, WA 99164; ALEXANDER MITRUSHCHENKOV, Laboratoire de Modélisation et Simulation Multi Echelle, Université Paris-Est Marne-la-Vallée, 77454 Marne la Vallée, Cedex 2, France; and JOSEPH S. FRANCISCO, Department of Chemistry, Purdue University, West Lafayette, IN 47907.

Over the last several years, composite methods have found great utility in the area of accurate ab initio thermochemistry. Utilizing highly correlated wavefunction-based methods such as CCSD(T) in conjunction with basis set extrapolations and corrections due to relativistic effects, core electron correlation, etc., accuracies approaching 1 kJ/mol have been possible in some cases. In the present work a similar methodology, including the use of explicitly correlated F12 methods and the inclusion of spin-orbit coupling, has been employed for the near-equilibrium potential energy surfaces of the ^{2}Π ground states of CCN and CCSb. A detailed analysis of the anharmonic vibrational spectra calculated from these surfaces and the Renner-Teller effect in these molecules will be discussed. The vibrational spectrum of the HNNO radical is found to be a challenging case for ab initio methods due to strong higher level electron correlation effects.^{a}

aK. A. Peterson and J. S. Francisco J. Chem. Phys. \underline{134}, 084308 (2011).